An arena receives $1,000 per event for 16 concerts. If the administrative costs for this sponsorship total $12,000, what is the profit margin for the concert sponsorship?

Answer by TimothyLamb(4379)

profit = 1000 * 16 - 12000 = 4000
profit margin = ( 4000 / (1000 * 16) ) * 100 = 25%
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php